# Memory Gym: Towards Endless Tasks to Benchmark Memory Capabilities of Agents

### v3

@@ Abstract
Keywords: deep reinforcement learning, proximal policy optimization, memory, transformer, recurrence
@@ 1 Introduction
Memory is not just a game – it is a critical tool for intelligent decision-making under imperfect information and uncertainty. Without the ability to recall past experiences, reasoning, creativity, planning, and learning may become elusive. In the realm of autonomously learning decision-making agents as Deep Reinforcement Learning (DRL), the agent’s memory involves maintaining a representation of previous observations, a knowledge bank that grounds its next decision. Memory mechanisms, be it through recurrent neural networks (Rumelhart et al., 1986) or transformers (Vaswani et al., 2017), have enabled these agents to master tasks both virtual and real. For instance, DRL methods have conquered complex video games as Capture the flag (Jaderberg et al., 2018), StarCraft \@slowromancapii@ (Vinyals et al., 2019), and DotA 2 (Berner et al., 2019). Their success extends beyond virtual environments to real-world challenges as dexterous in-hand manipulation (Andrychowicz et al., 2020) and controlling tokamak plasmas (Degrave et al., 2022).
@@ 1.1 Endless Tasks Benchmark Memory Effectiveness and not just Efficiency
Cumulative memory games, such as “I packed my bag”, introduce an automatic curriculum through their inherently progressive and unbounded nature. The game’s expansion depends on the participants’ continuous success. Exploiting the concept of ‘endlessness’ in tasks can provide deeper insights into the effectiveness of decision-makers when facing escalating demands of memory retention and recall. Such challenges inevitably test the limits of an agent’s memory. Here, ‘effectiveness’ refers to the ability to attain higher scores, indicative of more powerful memory capabilities. Contrasting this, traditional DRL memory benchmarks are centered around sample efficiency, influenced by the defined boundaries and fixed goals of finite tasks, which do not accommodate perpetual tasks.
@@ 1.2 Contributions: Novel Memory Benchmark and Transformer-XL Baseline
In our prior research, we introduced Memory Gym (Pleines et al., 2023), an open-source benchmark, designed to challenge memory-based DRL agents to memorize events across long sequences, generalize, be robust to noise, and be sample efficient. Memory Gym encompasses three distinct finite environments: Mortar Mayhem, Mystery Path, and Searing Spotlights. Each environment presents visual observations, multi-discrete action spaces, and crucially, they cannot be mastered without memory, demanding frequent memory interactions. Building on this foundation, we advance Memory Gym by evolving each environment into an endless format, inspired by the “I packed my bag” game, to further test and develop memory capabilities in DRL agents. To our current understanding, no existing DRL memory benchmark possesses such endless tasks.
We further contribute an open-source and easy-to-follow baseline implementation of Transformer-XL (TrXL) (Dai et al., 2019) based on the widely-used DRL algorithm Proximal Policy Optimization (PPO) (Schulman et al., 2017). A recurring theme in the realm of memory-based DRL is the lack of accessible and reproducible baselines. Many significant works, particularly those utilizing transformers (Fortunato et al., 2019; Parisotto et al., 2020; Hill et al., 2021; Lampinen et al., 2021; Parisotto and Salakhutdinov, 2021), have not made their implementations publicly available. As implementation details of DRL algorithms are vital (Engstrom et al., 2020; Andrychowicz et al., 2021; Huang et al., 2022b), this lack of transparency poses significant challenges for the research community, hindering progress and reproducibility. Therefore, our accessible implementation can serve as a beacon for the community fostering further research.
@@ 2 Related Work
Previous studies have explored the use of memory-based agents in a variety of tasks. Some of them are originally fully observable but are turned into partially observable Markov Decision Processes by adding noise or masking out information from the agent’s observation space (Hausknecht and Stone, 2015; Heess et al., 2015; Meng et al., 2021; Shang et al., 2021). Subsequently we give a coarse overview of recently used benchmarks chronologically.
@@
Miniworld (Chevalier-Boisvert, 2018) encompasses a set of first-person 3D environments based on the tasks introduced by Minigrid.
VizDoom (Wydmuch et al., 2018) features first-person shooter 3D environments that revolve around motor-control navigation tasks, with the added challenge of computer-controlled enemies that may confront the agent.
@@
Numpad (Parisotto et al., 2020) requires the agent to accurately press a series of keys in a specific sequence of fixed length. The sequence itself is not provided to the agent, resulting in the agent employing a trial-and-error approach while memorizing the underlying sequence.
Memory Maze (Parisotto et al., 2020) shares similarities with the Morris Water maze (D’Hooge and De Deyn, 2001), as the agent must locate an apple, then undergoes random repositioning, requiring it to find the apple’s location again and again.
@@
Memory Maze (Pašukonis et al., 2023) showcases procedurally generated 3D mazes, where agents, from a first-person perspective, are tasked with locating specific objects as requested by the environment. Additionally, the authors have collected a dataset tailored for offline reinforcement learning.
Memory Gym’s environments offer a unique feature that sets them apart from the aforementioned benchmarks: an endless design tailored to rigorously test memory capabilities. It is important to clarify that ‘endless’ should not be confused with ‘open-ended’. Memory Gym’s endless environments are not open-ended in nature; they possess clear objectives and defined structures. In contrast, open-ended environments offer agents freedom to develop unique skills and strategies without stringent constraints as demonstrated by the Paired Open-Ended Trailblazer algorithm (Wang et al., 2020) or platforms like Minecraft (Fan et al., 2022).
@@ 3 Endless Memory Tasks Inspired by “I Packed My Bag”
Prior to the underlying work, we introduced Memory Gym, a set of finite tasks benchmarking agents to strongly depend on frequent memory interactions (Pleines et al., 2023). Memory Gym consists of the environments Mortar Mayhem, Mystery Path, and Searing Spotlights. All of them expose visual observations, capturing an area of, $84\times 84$ RGB pixels, to the agent and allow for multi-discrete actions. Importantly, the original environments pose a single (or finite) goal. In this section, we delve into the intricacies of extrapolating the originally finite tasks to endless ones. Annex A presents the reset parameters to configure and scale the environments.
3.1 Endless Mortar Mayhem
@@
3.2 Endless Mystery Path
@@
3.3 Endless Searing Spotlights
@@ 4 Transformer-XL Baseline
In our experiments, we employ the widely-recognized Deep Reinforcement Learning (DRL) algorithm Proximal Policy Optimization (PPO) (Schulman et al., 2017). To add memory abilities to PPO, either recurrent neural networks (RNN) or attention mechanisms (i.e. transformer) are potential solutions. Prior works have extensively demonstrated the effectiveness of leveraging RNNs, such as Long Short-Term Memory (LSTM) (Hochreiter and Schmidhuber, 1997) and Gated Recurrent Unit (GRU) (Cho et al., 2014), within the realm of DRL (Mnih et al., 2016; Espeholt et al., 2018; Huang et al., 2022a). This also applies to previous studies employing transformers, as Gated Transformer-XL (GTrXL) (Parisotto et al., 2020) and Hierarchical Chunk Attention Memory (HCAM) (Lampinen et al., 2021). However, integrating these memory-enhancing techniques into DRL algorithms is not straightforward, particularly due to the dearth of readily available transformer-based DRL implementations. To bridge this gap, we contribute an easy-to-follow baseline implementation of Transformer-XL (TrXL) (Dai et al., 2019) in PPO. Before delving into the transformer-based PPO baseline, we describe the actor-critic model architecture and the losses used during training.
@@ 4.1 Actor-Critic Model Architecture
Optionally, there is a head dedicated to reconstructing the original visual observation to strengthen the learning signal: the latent output of the memory encoder is fed to a transposed Atari CNN. The utility of such an auxiliary head is demonstrated by some of our results (Section 5) and prior works by Lampinen et al. (2021) and Hill et al. (2021). During development, we also tried to attach the reconstruction head directly to the encoder, but this led to catastrophic results. Another optional head establishes a diagnosis tool by estimating ground truth information, which is provided as labels by the environment. Inspired by Baker et al. (2020), the parameters of this head can be optimized post agent training to assess whether distinct information can be queried from the agent’s memory.
@@ 4.2 Loss Function Composition
$L^{VClip}_{t}(\theta)=\left(\text{clip}\left(V(o_{t},h_{t},\theta),V(o_{t},h_{t},\theta_{\text{old}})-\epsilon,V(o_{t},h_{t},\theta_{\text{old}})+\epsilon\right)-\hat{V}\right)^{2}$ (2)
$L_{t}^{V}(\theta)=\max\left[\left(V(o_{t},h_{t},\theta)-\hat{V}\right)^{2},~{}L^{VClip}_{t}(\theta)\right]$ (3)
@@ 4.3 Transformer-XL as Memory Mechanism
(b) Transformer-XL Block
Figure 5: Figure 5(a) illustrates the episodic memory that stores past inputs to all Transformer-XL blocks for every step taken by the agent. Input sequences to the Transformer-XL blocks are retrieved using a sliding memory window. $T$ denotes the episode length. Figure 5(b) depicts the architecture of one Transformer-XL block that is stacked $L$ times. The dashed lines state that there is no gradient flow into the episodic memory, the absolute positional encoding, and the mask.
@@
To begin with, it needs to be clarified what kind of sequential data is fed to a transformer, which is typically composed of multiple stacked blocks. Naively, the input data can be conceptualized as a sequence of observations. In the context of visual observations, this is commonly referred to as frame stacking, where consecutive frames are combined. However, instead of expensively stacking raw frames, we leverage the past outputs of the model’s embedding module (Figure 4), which serves as the input sequence to the first transformer block. The subsequent blocks make use of the past outputs from their preceding blocks. These sequences, those elements we refer to as hidden states, are stored in the agent’s episodic memory (Figure 5(a)) and are collected during inference based on the episode’s current timestep $t$. To accommodate computational challenges of potentially endless episodes, we employ a sliding memory window approach, where the input sequence of past hidden states is derived from a fixed-length window. By leveraging formerly computed hidden states in the input sequence, the agent’s memory facilitates the ability to attend to events that extend beyond the boundaries of the fixed-length memory window, aligning with the segment-level recurrence of TrXL (Dai et al., 2019). Note that gradients are not back propagated through the episodic memory. Once propagated through all transformer blocks, a final hidden state is produced to inform the agent’s policy.
@@
To restrict attention to time steps up to the current timestep $t$, we employ a strictly lower triangular matrix as a mask in MHA. This mask is only necessary when $t$ is smaller than the length of the sliding memory window. The remaining configuration of the block adheres to the findings of Parisotto et al. (2020), which suggest improved performance with pre layer normalization and the identity map reordering. We encountered vanishing gradient issues when applying layer normalization after MHA and the fully connected layer, which is referred to as post layer normalization. Although our implementation supports the gating mechanism of GTrXL, it resulted in either lower or equivalent performance while being computationally more expensive (Appendix C).

### v5

@@ Abstract
Keywords: deep reinforcement learning, actor-critic, memory, transformer, recurrence
@@ 1 Introduction
Memory is not just a game – it is a critical tool for intelligent decision-making under imperfect information and uncertainty. Without the ability to recall past experiences, reasoning, creativity, planning, and learning may become elusive. In the realm of autonomously learning decision-making agents as Deep Reinforcement Learning (DRL), the agent’s memory involves maintaining a representation of previous observations, a knowledge bank that grounds its next decision. Memory mechanisms, be it through recurrent neural networks (Rumelhart et al., 1986) or transformers (Vaswani et al., 2017), have enabled these agents to master tasks both virtual and real. For instance, DRL methods have conquered complex video games as Capture the flag (Jaderberg et al., 2018), StarCraft \@slowromancapii@ (Vinyals et al., 2019), and DotA 2 (Berner et al., 2019). Their success extends beyond virtual environments to real-world challenges as dexterous in-hand manipulation (Andrychowicz et al., 2020) and controlling tokamak plasmas (Degrave et al., 2022). These non-Markovian examples underscore that effective memory mechanisms are crucial; without them, the tasks are fundamentally unsolvable as they require the robust maintenance and manipulation of information over time.
@@ 1.1 Endless Tasks Benchmark Memory Effectiveness and not just Efficiency
To discover the most effective memory approach for complex tasks as mentioned before, insightful benchmarks are needed. However, conventional DRL memory benchmarks, which typically revolve around finite tasks that terminate upon reaching a terminal state regardless of success or failure, fall short in providing the necessary depth. Finite tasks impose a predetermined limit on the demands placed on an agent’s memory. When various memory strategies successfully complete a fixed task, their performance is typically assessed based on efficiency rather than effectiveness. Here, “efficiency” refers to the number of agent-environment interactions required to achieve effective behaviors, commonly known as sample efficiency. Although these methods may reach the same end goal, their true effectiveness and memory capabilities often remain uncertain. For example, a memory mechanism that retains more information or sustains it over extended periods could prove more effective, potentially leading to higher scores, increased returns, better survival times, or improved win rates. This distinction could identify it as a more robust and effective approach compared to others assessed.
Cumulative memory games, such as “I packed my bag” and “Simon”(Morrison and Baer, 1977), overcome the limits of finite tasks. They introduce an automatic curriculum through their inherently progressive and unbounded nature. The game’s expansion depends on the participants’ continuous success. Exploiting the concept of “endlessness” in tasks provide deeper insights into the effectiveness of decision-makers when facing escalating demands of memory retention and recall. These challenges rigorously test an agent’s memory, evaluating both the quantity of information it can retain and the duration for which it can robustly maintain it.
@@ 1.2 Contributions: Novel Memory Benchmark and Transformer-XL Baseline
To exploit the endless behavior of cumulative memory games to thoroughly benchmark memory effectiveness, we enhance our prior work Memory Gym (Pleines et al., 2023), an open-source benchmark, designed to challenge memory-based DRL agents to memorize events across long sequences, generalize, be robust to noise, and be sample efficient. Memory Gym encompasses three distinct finite environments: Mortar Mayhem, Mystery Path, and Searing Spotlights. Each environment presents visual observations, multi-discrete action spaces, and crucially, they cannot be mastered without memory, demanding frequent memory interactions. Building on this foundation, we advance Memory Gym by evolving each environment into an endless format, inspired by the “I packed my bag” game, to further test and develop memory capabilities in DRL agents. standing, no existing DRL memory benchmark possesses such endless tasks.
We further contribute an open-source and easy-to-follow baseline implementation of Transformer-XL (TrXL) (Dai et al., 2019) based on the widely-used DRL algorithm Proximal Policy Optimization (PPO) (Schulman et al., 2017). A recurring theme in the realm of memory-based DRL is the lack of accessible and reproducible baselines. Many significant works, particularly those utilizing transformers (Fortunato et al., 2019; Parisotto et al., 2020; Hill et al., 2021; Lampinen et al., 2021; Parisotto and Salakhutdinov, 2021), have not made their implementations publicly available. As implementation details of DRL algorithms are vital (Engstrom et al., 2020; Andrychowicz et al., 2021; Huang et al., 2022), this lack of transparency poses significant challenges for the research community, hindering progress and reproducibility. Therefore, our accessible implementation can serve as a beacon for the community fostering further research.
@@ 2 Related Work
Previous studies have explored the use of memory-based agents in a variety of tasks. Some of them are originally fully observable but are turned into partially observable Markov Decision Processes by adding noise or masking out information from the agent’s observation space (Hausknecht and Stone, 2015; Heess et al., 2015; Meng et al., 2021; Shang et al., 2021). In the next subsection, we give a coarse overview of recently used benchmarks chronologically. Subsequently, we discuss how these benchmarks primarily assess efficiency, highlighting a limitation in their scope, which we address with our contribution of Memory Gym’s endless environments.
2.1 An Overview of Memory Benchmarks in Deep Reinforcement Learning
@@
Miniworld (Chevalier-Boisvert, 2018) encompasses a set of first-person 3D environments sharing similarities to the tasks introduced by Minigrid.
VizDoom (Wydmuch et al., 2019) features first-person shooter 3D environments that revolve around motor-control navigation tasks, with the added challenge of computer-controlled enemies that may confront the agent.
@@
Numpad (Parisotto et al., 2020) requires the agent to accurately press a series of keys in a specific sequence of fixed length. The sequence is not provided to the agent, resulting in the agent using a trial-and-error approach while memorizing the underlying sequence.
Memory Maze (1) (Parisotto et al., 2020) shares similarities with the Morris Water maze (D’Hooge and De Deyn, 2001), as the agent must find an apple, then undergoes random repositioning, requiring it to relocate the apple’s location.
@@
Memory Maze (2) (Pašukonis et al., 2023) showcases procedurally generated 3D mazes, where agents, from a first-person perspective, are tasked with locating specific objects as requested by the environment. Additionally, the authors have collected a dataset tailored for offline reinforcement learning.
2.2 Conventional Benchmarks are Limited to Measuring Efficiency
All of the aforementioned environments conclude with terminal states, regardless of whether the agent succeeds or fails. Navigational tasks such as those found in Deepmind Lab 30, Minigrid, Miniworld, VizDoom, Procgen, Numpad, Memory Maze (1), and Ballet end by the agent discovering an exit. This closure also applies to the navigation challenges within the Memory Task Suite and PopGym. Additional tasks in PopGym conclude after the agent plays through an entire deck of cards or reaches a distinct goal state in games such as Concentration, Battleship, and Minesweeper. Memory Maze (2) and tasks in Psychlab terminate after a predefined number of steps.
Due to these clearly finite episodes, the demands on the agent’s memory have an upper bound. Once the agent solves a given task, their performance is primarily measured by their efficiency. Efficiency here refers to the number of steps an agent requires to complete an episode. In scenarios such as escaping a maze, a more efficient agent will use fewer steps, whereas even a random agent might eventually find the exit if given sufficient time. In both the literature and our results on the finite environments of Memory Gym (Section 5.3), agents are typically compared based on the number of samples (i.e., agent-environment interactions) they use during training to achieve their performance. But what if an agent’s memory is stronger at recalling larger amounts of relevant information over extended horizons? Efficiency metrics may fail to fully reveal the true memory capabilities of agents, including their capacity to retain larger amounts of information or sustain that information over extended periods.
While the related works still offer potential for scaling in terms of memory demands, we take inspiration from cumulative memory games to expand Memory Gym to provide endless tasks that automatically and boundlessly test memory capabilities. Identifying particularly strong memory mechanisms may enable the tackling of complex tasks, as noted in the introduction.
@@ 3 Endless Memory Tasks Inspired by “I Packed My Bag”
Memory Gym’s environments offer a unique feature that sets them apart from the related benchmarks: an endless design tailored to rigorously test memory capabilities. This section first explores how exploiting the concept of cumulative memory games can serve as a valuable benchmark to assess memory effectiveness, not merely efficiency. We then describe how we expand Memory Gym’s finite environments into endless ones.
It is important to clarify that ‘endless’ should not be confused with ‘open-ended’. Memory Gym’s endless environments are not open-ended in nature; they possess clear objectives and defined structures. In contrast, open-ended environments offer agents freedom to develop unique skills and strategies without stringent constraints as demonstrated by the Paired Open-Ended Trailblazer algorithm (Wang et al., 2020) or platforms like Minecraft (Fan et al., 2022).
3.1 Evaluating Effectiveness through Cumulative Memory Games
As argued in Section 2.2, measuring sample efficiency may obscure the true effectiveness of an agent’s memory capabilities, especially under the assumption that the evaluated baselines are sufficient to solve the given finite tasks. If endless tasks are not considered, what alternative methods can be used to assess an agent’s memory effectiveness? There are two primary approaches: scaling the difficulty of the task or scaling the agent.
One approach involves progressively ablating the agent itself while maintaining a fixed task difficulty. For instance, the dimensions of transformer architectures or recurrent cells are scalable in terms of their parameter count, which is indicative of the memory’s capacity. By iteratively reducing these dimensions, a threshold is eventually reached beyond which the agent can no longer perform effectively. Thus, with a fixed task, we can only determine the minimal configuration that still allows the agent to be effective.
In contrast, incrementally increasing the task’s difficulty can determine the point at which an agent fails, thereby exposing the agent’s limits of effectiveness. The field of Curriculum Learning (CL), a specialization of transfer learning, offers methods to automate this scaling process (Narvekar et al., 2020). CL enhances training efficiency by sequentially transferring knowledge from simpler tasks to more challenging ones, ultimately addressing tasks that might be too complex to tackle from scratch. While setting up curricula could introduce implementation overhead, such as hand-crafting tasks or scheduling transitions between tasks, cumulative memory games naturally function out-of-the-box as an automatic curriculum. These endless games increase in difficulty as the agent progresses within an episode, using a trivial heuristic to adjust the challenge in an online fashion, thereby eliminating the need for a separate task scheduler.
Applying the concept of cumulative memory games to a DRL memory benchmark provides significant insights. Agents that excel at these tasks demonstrate robustness in retaining and recalling vast amounts of information over extended periods. Moreover, because these games offer unbounded difficulty, they remain appropriately challenging—neither too difficult nor too easy—and hence continue to provide the right challenge for evaluating future memory mechanisms. Mastery of these memory-intensive games highlights an agent’s suitability for real-world applications, where maintaining context and adapting to incremental information are essential.
3.2 From Finite to Endless: Expanding Memory Gym’s Environments
Prior to the underlying work, we introduced Memory Gym, a set of finite tasks benchmarking DRL agents to strongly depend on frequent memory interactions (Pleines et al., 2023). Memory Gym consists of the environments Mortar Mayhem, Mystery Path, and Searing Spotlights. All of them expose visual observations, capturing an area of $84\times 84$ RGB pixels, to the agent and allow for multi-discrete actions. Leveraging procedural content generation, each episode is restarted with a new level sampled from a distinct set of seeds. Additionally, the environments’ simulation speed measures multiple thousands of steps per second, which is faster than most related benchmarks (see Appendix F). Originally, these tasks were designed with finite goals, which limited the scope of memory challenges presented to the agents.
In transforming these environments from finite to endless, we aim to enhance the challenge posed to an agent’s memory by increasing the amount of information to memorize over expanding time:
  • Endless Mortar Mayhem tasks the agent to memorize and execute an ever-expanding list of commands in the correct order.
  • Endless Mystery Path exhibits a never-ending, invisible path for the agent to traverse. Deviating from the path forces a restart from the beginning, requiring the agent to recall both the route and previous missteps accurately.
  • Endless Searing Spotlights demands the agent to track its obscured location while avoiding threatening spotlights and collecting coins.
For detailed configuration and scaling of these environments, refer to Annex A, which presents the reset parameters. The next subsections explore the transitions from finite tasks to endless ones.
3.2.1 Endless Mortar Mayhem
@@
3.2.2 Endless Mystery Path
@@
3.2.3 Endless Searing Spotlights
@@ 4 Transformer-XL Baseline
In our experiments, we employ the widely-recognized Deep Reinforcement Learning (DRL) algorithm Proximal Policy Optimization (PPO) (Schulman et al., 2017). To add memory abilities to PPO, either recurrent neural networks (RNN) or attention mechanisms (i.e. transformer) are potential solutions. Prior works have extensively demonstrated the effectiveness of leveraging RNNs, such as Long Short-Term Memory (LSTM) (Hochreiter and Schmidhuber, 1997) and Gated Recurrent Unit (GRU) (Cho et al., 2014), within the realm of DRL (Mnih et al., 2016; Espeholt et al., 2018; Huang et al., 2022). This also applies to previous studies employing transformers, as Gated Transformer-XL (GTrXL) (Parisotto et al., 2020) and Hierarchical Chunk Attention Memory (HCAM) (Lampinen et al., 2021). However, integrating these memory-enhancing techniques into DRL algorithms is not straightforward, particularly due to unavailable transformer-based DRL implementations. To bridge this gap, we contribute an easy-to-follow baseline implementation of Transformer-XL (TrXL) (Dai et al., 2019) in PPO. Before delving into the transformer-based PPO baseline, we describe the actor-critic model architecture and the losses used during training.
@@ 4.1 Actor-Critic Model Architecture
Optionally, there is a head dedicated to reconstructing the original visual observation to strengthen the learning signal: the latent output of the memory encoder is fed to a transposed Atari CNN. The utility of such an auxiliary head is demonstrated by some of our results (Section 5) and prior works by Lampinen et al. (2021) and Hill et al. (2021). During development, we also tried to attach the reconstruction head directly to the encoder, but this led to catastrophic results. Another optional head establishes a diagnosis tool by estimating ground truth information, which is provided as labels by the environment. Despite the injection of illegal information into the agent’s training, we utilize this head in Section 5.6.2 to demonstrate that TrXL’s performance improves in Endless Mortar Mayhem due to this richer learning signal, indicating that its overall low performance is not attributed solely to a lack of capacity. Initially, the parameters of such a head can be optimized post-agent training to evaluate whether distinct information can be queried from the agent’s memory (Baker et al., 2020).
@@ 4.2 Loss Function Composition
$L^{VClip}_{t}(\theta)=\left(\text{clip}\left(V(o_{t},h_{t},\theta),V(o_{t},h_{t},\theta_{\text{old}})-\epsilon,V(o_{t},h_{t},\theta_{\text{old}})+\epsilon\right)-\hat{V}_{t}\right)^{2}$ (2)
$L_{t}^{V}(\theta)=\max\left[\left(V(o_{t},h_{t},\theta)-\hat{V}_{t}\right)^{2},~{}L^{VClip}_{t}(\theta)\right]$ (3)
@@ 4.3 Transformer-XL as Memory Mechanism
(b) Transformer-XL Layer
Figure 5: Figure 5(a) illustrates the episodic memory that stores past inputs to all Transformer-XL layers for every step taken by the agent. Input sequences to the Transformer-XL layers are retrieved using a sliding memory window. $T$ denotes the episode length. Figure 5(b) depicts the architecture of the Transformer-XL block that stacks $L$ layers. The dashed lines state that there is no gradient flow into the episodic memory, the absolute positional encoding, and the mask.
@@
To begin with, it needs to be clarified what kind of sequential data is fed to a transformer, which is typically composed of multiple stacked layers. Naively, the input data can be conceptualized as a sequence of observations. In the context of visual observations, this is commonly referred to as frame stacking, where consecutive frames are combined. However, instead of expensively stacking raw frames, we leverage the past (i.e. cached) outputs of the model’s embedding module (Figure 4), which serves as the input sequence to the first transformer layer. The subsequent layers make use of the past outputs from their preceding layers. These sequences, those elements we refer to as cached hidden states, are stored in the agent’s episodic memory (Figure 5(a)) and are collected during inference based on the episode’s current timestep $t$. To accommodate computational challenges of potentially endless episodes, we employ a sliding memory window approach, where the input sequence of these hidden states is derived from a fixed-length window. By leveraging cached hidden states in the input sequence, the agent’s memory facilitates the ability to attend to events that extend beyond the boundaries of the fixed-length memory window, aligning with the segment-level recurrence of TrXL (Dai et al., 2019). Note that gradients are not back propagated through the episodic memory because of caching. Once propagated through all transformer layers, a final hidden state is produced to inform the agent’s policy.
@@
To restrict attention to time steps up to the current timestep $t$, we employ a strictly lower triangular matrix as a mask in MHA. This mask is only necessary when $t$ is smaller than the length of the sliding memory window. The remaining configuration of the block adheres to the findings of Parisotto et al. (2020), which suggest improved performance with pre layer normalization and the identity map reordering. We encountered vanishing gradient issues when applying layer normalization after MHA and the fully connected layer, which is referred to as post layer normalization. Although our implementation supports the gating mechanism of GTrXL, it resulted in either lower or equivalent performance while being computationally more expensive (Appendix D).